OpenStudy (anonymous):

how to solve this equation? 5x+4y+6z=5 2x-4y-6z=2 2x+3y+x=9 solve for the x,y and z

OpenStudy (bahrom7893):

have you taken linear algebra yet?

OpenStudy (bahrom7893):

I guess not.. oh this is going to be something very long without matrices..

OpenStudy (anonymous):

our lesson is about Solving Systems of Linear Equations by Elimination

OpenStudy (bahrom7893):

Okay let's eliminate one variable. Suppose you solve for 4y in the top equation: 4y = 5 - 6z - 5x Plug that into the 2nd equation. What do u get?

OpenStudy (bahrom7893):

2x-4y-6z=2 2x - (5 - 6z - 5x) -6z = 2

OpenStudy (bahrom7893):

2x - 5 + 6z + 5x - 6z = 2 2x + 5x -5 = 2 7x = 7 x =1

OpenStudy (bahrom7893):

2x+3y+x=9 2*1 + 3y + 1 = 9 2 + 1 +3y = 9 3 + 3y = 9 3y = 6 y = 2

OpenStudy (moongazer):

We have the same lesson

OpenStudy (bahrom7893):

4y = 5 - 6z - 5x 4*2 = 5 - 6z - 5*1 8 = -6z z = -8/6 z = -4/3 just check my arithmetic..

OpenStudy (moongazer):

Are you in fourth year high school?

OpenStudy (anonymous):

nope...stll in 2nd year:))

OpenStudy (moongazer):

We had a quiz about this little a while ago and i got ZERO even if I know how to do it I just don't know why I can't solve it. I had a perfect score in my seatwork but in my QUIZ it is ZERO what do you think is my problem. my quiz is over 20

OpenStudy (moongazer):

bahrom your method is different from my teacher taught us

OpenStudy (bahrom7893):

moongazer then use yours and explain it to colleen.. Ive always used this..

OpenStudy (moongazer):

could you teach me?

OpenStudy (bahrom7893):

teach u what? my method?

OpenStudy (moongazer):

because I think yours is easier

OpenStudy (moongazer):

yup

OpenStudy (moongazer):

I will give a sample

OpenStudy (moongazer):

i will make a question about this

OpenStudy (bahrom7893):

okay give me an example.. preferably 3 eqns and 3 unknowns.. but works with any # of eqns and unknowns if they are equal to each other that is..